1) The price of spring water rises from $1.90 to $2.10 a bottle, and the quantity demanded decreases from 11 million to 9 million bottles a week.
a) Calculate the percentage change in the price of spring water.
b) Calculate the percentage change in the quantity demanded of spring water.
c) Is the demand for spring water elastic or inelastic?
d) Would the demand for Pepsi be more elastic or less elastic than the demand for spring water? Why?
e) Calculate the price elasticity of demand for spring water.
f) At what price is the price elasticity of demand for spring water equal to your answer in part (e)?
g) What is the change in the total revenue of sellers of spring water?
h) If the demand curve for spring water is a straight-line demand curve, is the price at which the demand for spring water is unit elastic a higher price or a lower price than your answer to part (f)? Why
1
Expert's answer
2014-11-28T14:47:27-0500
1) The price of spring water rises from $1.90 to $2.10 a bottle, and the quantity demanded decreases from 11 million to 9 million bottles a week.
a) The percentage change in the price of spring water is (2.1/1.9 - 1)*100% = 10.5%
b) The percentage change in the quantity demanded of spring water is (9/11 - 1)*100% = -18.2%
c) The demand for spring water is elastic.
d) The demand for Pepsi would be less elastic than the demand for spring water, because Pepsi is more differentiated product.
e) The price elasticity of demand for spring water Ed = -18.2/10.5 = -1.73, |Ed| > 1.
f) The price elasticity of demand for spring water is equal to the answer in part at price of (2.1 - 1.9)/2 = $2
g) The change in the total revenue of sellers of spring water is 2.1*9 - 1.9*11 = -$2 million
h) If the demand curve for spring water is a straight-line demand curve, the price at which the demand for spring water is unit elastic is a lower price than in the answer to part (f).